(d) The following is an extract from the national income data of a country in 2020. Study the extract and answer the question(d) The following is an extract from the national income data of a country in 2020. Study the extract and answer the questions that follow: The agricultural sector in 2020 produced foodstuffs, livestock, coffee and groundnuts worth GHȼ 1,151 million. Manufacturing, construction, mining, water and electricity produced GHȼ 1,035 million. Transport and communication, finance, insurance and government services contributed GHȼ 1,170 million. Farmers received GHȼ 350 million as subsidies while government collected GHȼ 390 million in indirect taxes. GHC150 million worth of capital was estimated to have been used up. Required: i. Determined the country’s Gross Domestic Product in market prices. (4 marks) ii. Calculate the Net Domestic Product at factor cost. (4 marks) iii. What method of estimating the national income is used in the extract? (2 marks)
Answer #1

d)i) The GDP in market prices equals the total value of goods produced by different sectors in the nation - subsidies + indirect taxes.

Total value of goods poduced = Agricultural sector-GHȼ 1,151 million + Manufacturing, construction, mining, water and electricity-GHȼ 1,035 million + Transport and communication, finance, insurance and government services-GHȼ 1,170 million = GHȼ 3,356 million.

GDP at MP = GHȼ (3,356 - 350 + 390)million = GHȼ 3,396 million

ii) NDP at FC = GDP at MP - Indirect taxes + subsidy - depreciation

= GHȼ (3,396 - 390 + 350 - 150)million = GHȼ 3,206 million

iii) In the given extract we use the product method to give the value of GDP. It gives the total value of goods produced by each sector of the nation